OpenStudy (anonymous):

Given the area of a triangle is 40 cm2, if its altitude is 8 cm, its base is 15 cm. True False

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

Im thinking its false but not sure

OpenStudy (anonymous):

8x15=120

OpenStudy (mayankdevnani):

area of triangle=1/2*base*height

OpenStudy (mayankdevnani):

1/2*8*15=60cm^2

OpenStudy (mayankdevnani):

yaa it is false

OpenStudy (mayankdevnani):

@Claire4christ ok

OpenStudy (anonymous):

A rectangle has an area of 150 m². If one side is 25 m, then the length of the adjacent side is how long? 60 m 37.5 m 0.6 m 6 m

OpenStudy (anonymous):

6

OpenStudy (mayankdevnani):

area of rectangle=l*b 150=25*x x=6

OpenStudy (mayankdevnani):

option d

OpenStudy (anonymous):

ok i got it

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

A square measures 6 inches on each side. What is it's area rounded off to the nearest whole number. 232 cm² 1,296 in.² 91 cm² 18 in.²

OpenStudy (anonymous):

little confused

OpenStudy (anonymous):

1=2.5 6x4=24 24x2.5=60

OpenStudy (anonymous):

false . 40 = 1/2 b(8) b=10 .

OpenStudy (mayankdevnani):

area of square=side^2=6^2=36

OpenStudy (mayankdevnani):

36cm^2 isn't options

OpenStudy (anonymous):

what

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

comvert into inches

OpenStudy (anonymous):

explain

OpenStudy (mayankdevnani):

1 inch equals to

OpenStudy (anonymous):

i have to get one of thoses answers above

OpenStudy (anonymous):

2.5

OpenStudy (mayankdevnani):

so multiply 2.5 by 36=

OpenStudy (anonymous):

90